In 2013, there were 19,767 Starbucks stores world- wide. This is approximately 4551% more stores than in 1994. How many Starbucks stores were there in 1994 ? (Round to the nearest whole number.)
1
Expert's answer
2017-10-20T10:39:06-0400
Let x – number of StarBucks in 1994
X = 100%
X + 4551%(x) = 100% + 4551% = 4651%
4651% = 19767
1% = 19767 / 4651 = 4.25
X = 4.25*100% = 425
Answer : 425 Starbucks In 1994